# On the Real Zeroes of Half-integral Weight Hecke Cusp Forms, II

### v1

@@ 10.3. Estimates for primes $p>K^{\theta_{J}}$
$\ll XK(\log K)^{2}\sum_{n\leq K^{4\theta_{J}}}\frac{\widetilde{h}_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\sum_{\begin{subarray}{c}u|n\\ nu=\square\end{subarray}}\frac{c_{n}(u)d(u)}{\sqrt{u}}\prod_{p|n}\left(1+\frac{1}{p}\right)^{-1}$
@@
We remove the condition $\Omega(n)\leq 4\ell_{j}$ by arguing as before: using Rankin’s trick and the estimate $c_{n}(u)\leq 2^{\Omega(n)}$ it is easy to see that there exists $C>0$ such that
@@
Several positive constants appear throughout the proof. We start by summarising their roles:
@@
    $C_{1}$ we are able to choose freely.
@@
    $C_{2}$ we are able to choose freely.
@@
    $C_{5}$ we are able to choose freely.
@@
    $C_{7}$ we are able to choose freely.
@@
    $C_{11}$ we are able to choose freely.
@@
    $C_{12}:=4C_{11}$.
We require these to satisfy following constraints:
@@
Note that the existence of the constants $C_{3}$ and $C_{4}$ follows from Propositions 4.1 and 4.2, respectively. Recall that the constants $C_{3}$, $C_{4}$ and $C_{9}$ are absolute. The first condition is satisfied by choosing $C_{5}$ to be large enough. Similarly, if we have chosen $C_{8}$ so that it is large enough, the second condition holds by choosing $C_{7}$ to be small enough. The third condition is true when $C_{11}$ is chosen to be small enough. When $C_{11}$ is chosen, the fourth condition holds when $C_{2}$ is chosen suitably. The final condition is met, note that $C_{5}$, $C_{3}$, $C_{7}$, $C_{11}$ and $C_{2}$ are already fixed, by choosing $C_{1}$ to be large enough.
For the rest of the paper, set $X=K/Y$. Then $K^{\vartheta}\ll X\ll\sqrt{K/\log K}$ by the assumptions on $Y$. Remember that in order to detect sign changes along the sequence $d\equiv 1\,(\text{mod }4)$ with $d$ squarefree and $(-1)^{k}d>0$ on the short interval $[x,x+H]$, $x\sim X$, it suffices to have
@@
Choose $h$ to be a non-negative smooth function that is supported in the interval $[1/2,5/2]$ and is identically one in $[1,2]$. With the above notation the quantity we need to bound is by Markov’s inequality
@@
$=\frac{1}{C_{2}X}\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\sum_{g\in B_{k+\frac{1}{2}}^{+}}\sum_{\begin{subarray}{c}x\sim X\\ |S_{1,g}(x;H)|\geq\sqrt{C}_{1}\cdot Hk^{-1/2}\end{subarray}}1$
@@
The off-diagonal contribution is negligible identically as in the proof of [13, Proposition 2.5.].
@@
$\mathcal{V}_{g}:=\left\{x\sim X:\,\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ M_{g}(d)^{2}L(1/2,f\otimes\chi_{d})\leq C_{5}\end{subarray}}\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)\geq\frac{C_{7}H}{k^{1/2}}\right\}.$
@@
$=\frac{C_{8}}{H}\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\sum_{g\in B_{k+\frac{1}{2}}^{+}}\left(\sum_{x\in\mathcal{V}_{g}}+\sum_{x\not\in\mathcal{V}_{g}}\right)\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ M_{g}(d)^{2}L(1/2,f\otimes\chi_{d})\leq C_{5}\end{subarray}}\alpha_{g}^{1/2}\omega_{g}^{1/2}|c_{g}(|d|)|$
$\leq\frac{C_{8}}{H}\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\sum_{g\in B_{k+\frac{1}{2}}^{+}}\left(C_{5}\omega_{g}H\left|\mathcal{V}_{g}\right|+\omega_{g}^{1/2}\frac{C_{7}XH}{k^{1/2}}\right),$
@@
$\mathcal{U}:=\left\{g\in\mathcal{S}_{K}:\,|\mathcal{V}_{g}|>C_{11}X\right\}.$
@@
by recalling that $C_{12}=4C_{11}$ and $C_{11}$ is chosen so that $C_{12}<C_{10}$.
@@
$\#\left\{x\sim X\,:\,\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{x\leq(-1)^{k}d\leq x+H}\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)\geq\frac{C_{7}H}{k^{1/2}}\right\}$
$\geq\#\left\{x\sim X\,:\,\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ M_{g}(d)^{2}L(1/2,f\otimes\chi_{d})\leq C_{5}\end{subarray}}\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)\geq\frac{C_{7}H}{k^{1/2}}\right\}$
$>C_{11}X.$
@@
for almost all $x\sim X$ with the exceptional set having size $\leq(1-C_{2})X$. Moreover, for at least $(C_{10}-C_{12})K^{2}$ of the forms $g\in S_{K}$ it holds that
@@
holds for $\geq(C_{11}-1-C_{2})X$ of the numbers $x\sim X$ if we choose $H>(C_{1}/C_{7})^{2}$.
@@
holds for $\geq(C_{11}-1-C_{2})X$ of the numbers $x\sim X$. We note that the contribution coming from the summands with $\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)\leq k^{-\delta}$ is trivially bounded by, say,
$\leq\sum_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)\leq k^{-\delta}\end{subarray}}\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)\leq 2HK^{-\delta}$
as $\eta_{2}$ is sufficiently small, and so we conclude, choosing $\delta=1/2+\eta/2$ for concreteness so that $\delta>1/2$, that for the same proportion of $g\in\mathcal{S}_{K}$ and $x\sim X$ we have
$\left|\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)>k^{-\delta}\end{subarray}}\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)\,\pm\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)>k^{-\delta}\end{subarray}}\sqrt{\alpha_{g}}c_{g}(|d|)M_{g}(d)\right|\geq\frac{(C_{7}H-C_{1}\sqrt{H})}{\sqrt{k}}>0.$
Now observe that this implies
$2\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}c_{g}(|d|)M_{g}(d)>k^{-\delta}\end{subarray}}\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)$
$=\left|\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)>k^{-\delta}\end{subarray}}\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)\,+\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)>k^{-\delta}\end{subarray}}\sqrt{\alpha_{g}}c_{g}(|d|)M_{g}(d)\right|$
@@
$2\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}c_{g}(|d|)M_{g}(d)<-k^{-\delta}\end{subarray}}\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)$
$=\left|\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)>k^{-\delta}\end{subarray}}\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)\,-\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)>k^{-\delta}\end{subarray}}\sqrt{\alpha_{g}}c_{g}(|d|)M_{g}(d)\right|$
@@
Thus we have shown that for $\geq(C_{10}-C_{12}-2C_{3}/C_{2}C_{1}^{2})K^{2}$ of the forms $g\in\mathcal{S}_{K}$ the short interval $[x,x+H]$, $x\sim X$, contains numbers $(-1)^{k}d_{\pm}$, with both $d_{\pm}$ odd fundamental discriminants, for which $\sqrt{\alpha_{g}}c_{g}(|d_{+}|)M_{g}(d_{+})>k^{-\delta}$ and $\sqrt{\alpha_{g}}c_{g}(|d_{-}|)M_{g}(d_{-})<-k^{-\delta}$, for $\geq(C_{11}-1-C_{2})X$ of the numbers $x\sim X$ as $M_{g}(d)\ll(\log K)^{O(1)}$. Thus we deduce a sign change of $c_{g}(|d|)$ over a positive proportion of intervals of constant size $(C_{1}/C_{7})^{2}$ for a positive proportion of forms $g\in S_{K}$. This leads to

### v2

@@ 10.3. Estimates for the large primes
$\ll XK(\log K)^{2}\sum_{n\leq K^{4\delta_{0}}}\frac{\widetilde{h}_{2}(n)\lambda(n)}{2^{\Omega(n)}\sqrt{n}}\sum_{\begin{subarray}{c}u|n\\ nu=\square\end{subarray}}\frac{c_{n}(u)d(u)}{\sqrt{u}}\prod_{p|n}\left(1+\frac{1}{p}\right)^{-1}$
@@
We remove the condition $\Omega(n)\leq 4\ell_{j}$ by arguing as before: using Rankin’s trick and the estimate $c_{n}(u)\leq 2^{\Omega(n)}$ it is easy to see (as in the proof of (10.2)) that there exists $C>0$ such that
@@
Several positive constants appear throughout the proof. We start by summarising their roles. The order in which we fix the constants matters as the constants fixed later may depend on the constants chosen earlier in order to satisfy certain requirements. The constant $C_{5}$, $C_{8}$, $C_{7}$, $C_{2}$, $C_{11}$, and $C_{1}$ are fixed in this order.
@@
    $C_{1}$ we are able to choose freely provided that it is large enough compared to $C_{7}$.
@@
    $C_{2}$ we are able to choose freely provided that it is <1<1.
@@
    $C_{5}$ we are able to choose freely provided that it is sufficiently large.
@@
    $C_{7}$ we are able to choose freely provided that it is sufficiently small in terms of $C_{8}$.
@@
    $C_{11}$ we are able to choose freely provided that it is arge enough in terms of $C_{2}$.
@@
    $C_{12}:=C_{11}/4$.
We require these to satisfy the following constraints:
@@
  6. (6)
    $C_{1}>C_{7}$.
Note that the existence of the constants $C_{3}$ and $C_{4}$ follows from Propositions 4.1 and 4.2, respectively. Recall that the constants $C_{3}$, $C_{4}$ and $C_{9}$ are absolute. The first condition is satisfied by choosing $C_{5}$ to be large enough. Similarly, if we have chosen $C_{8}$ so that it is large enough, the second condition holds by choosing $C_{7}$ to be small enough. The third condition is true when $C_{11}$ is chosen to be small enough. When $C_{11}$ is chosen, the fourth condition holds when $C_{2}$ is chosen suitably. The fifth condition is met, note that $C_{5}$, $C_{3}$, $C_{7}$, $C_{11}$ and $C_{2}$ are already fixed, by choosing $C_{1}$ to be large enough. This choice of $C_{1}$ also satisfies the final condition if $C_{1}$ to be sufficiently large compared to $C_{7}$.
For the rest of the paper, set $X=K/Y$. Then $K^{\vartheta}\ll X\ll\sqrt{K/\log K}$ by the assumptions on $Y$. Remember that in order to detect sign changes along the sequence $d\equiv 1\,(\text{mod }4)$ with $d$ squarefree and $(-1)^{k}d>0$ in the short interval $[x,x+H]$, $x\sim X$, it suffices to have
@@
Choose $h$ to be a non-negative smooth function that is supported in the interval $[1,9/2]$ and is identically one in $[3/2,4]$. With the above notation the quantity we need to bound is by Markov’s inequality
@@
$=\frac{1}{C_{2}X}\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\sum_{g\in B_{k+\frac{1}{2}}^{+}}\sum_{\begin{subarray}{c}x\sim X\\ |S_{1,g}(x;H)|\geq\sqrt{C}_{1}Hk^{-1/2}\end{subarray}}1$
@@
The off-diagonal contribution is negligible identically as in the proof of [13, Proposition 2.5.] using the trivial bound $M_{g}(d)\ll K^{\delta_{0}}(\log K)^{1/4}$.
@@
$\mathcal{V}_{g}:=\left\{x\sim X:\,\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ M_{g}(d)^{2}L(1/2,f\otimes\chi_{d})\leq C_{5}\end{subarray}}\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)\geq\frac{C_{7}H}{\sqrt{k}}\right\}.$
@@
$=\frac{C_{8}}{H}\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\sum_{g\in B_{k+\frac{1}{2}}^{+}}\left(\sum_{x\in\mathcal{V}_{g}}+\sum_{x\not\in\mathcal{V}_{g}}\right)\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ M_{g}(d)^{2}L(1/2,f\otimes\chi_{d})\leq C_{5}\end{subarray}}\alpha_{g}^{1/2}\omega_{g}^{1/2}|c_{g}(|d|)|M_{g}(d)$
$\leq\frac{C_{8}}{H}\sum_{k\in\mathbb{Z}}h\left(\frac{2k-1}{K}\right)\sum_{g\in B_{k+\frac{1}{2}}^{+}}\left(\sqrt{C_{5}}\omega_{g}H\left|\mathcal{V}_{g}\right|+\omega_{g}^{1/2}\frac{C_{7}XH}{\sqrt{k}}\right),$
@@
$\mathcal{U}:=\left\{g\in\mathcal{S}_{K}:\,|\mathcal{V}_{g}|\geq C_{11}X\right\}.$
@@
by recalling that $C_{12}=C_{11}/4$ and $C_{11}$ is chosen so that $C_{12}<C_{10}$.
@@
$\#\left\{x\sim X\,:\,\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{x\leq(-1)^{k}d\leq x+H}\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)\geq\frac{C_{7}H}{\sqrt{k}}\right\}$
$\geq\#\left\{x\sim X\,:\,\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ M_{g}(d)^{2}L(1/2,f\otimes\chi_{d})\leq C_{5}\end{subarray}}\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)\geq\frac{C_{7}H}{\sqrt{k}}\right\}$
$\geq C_{11}X.$
@@
for a positive proportion of $x\sim X$ with the exceptional set having size $\leq(1-C_{2})X$. Moreover, for at least $(C_{10}-C_{12})K^{2}$ of the forms $g\in S_{K}$ it holds that
@@
holds for $\geq(C_{11}-1+C_{2})X$ of the numbers $x\sim X$ if we choose $H>(C_{1}/C_{7})^{2}$.
@@
holds for $\geq(C_{11}-1+C_{2})X$ of the numbers $x\sim X$. We note that the contribution coming from the summands with $\sqrt{\alpha_{g}}|c_{g}(|d|)\leq k^{-\delta}$ is trivially bounded by, say,
$\leq\sum_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}|c_{g}(|d|)|\leq k^{-\delta}\end{subarray}}\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)\leq 2K^{-\delta}\sum_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}|c_{g}(|d|)|\leq k^{-\delta}\end{subarray}}M_{g}(d)\leq 2HK^{-\delta+\delta_{0}}(\log K)^{1/4}$
as $\eta_{1}$ is sufficiently small, and so we conclude, choosing $\delta=1/2+\eta/2$ for concreteness so that $\delta>1/2$, that for the same proportion of $g\in\mathcal{S}_{K}$ and $x\sim X$ we have
$\left|\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}|c_{g}(|d|)|>k^{-\delta}\end{subarray}}\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)\,\pm\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}|c_{g}(|d|)|>k^{-\delta}\end{subarray}}\sqrt{\alpha_{g}}c_{g}(|d|)M_{g}(d)\right|\geq\frac{(C_{7}H-C_{1}\sqrt{H})}{\sqrt{k}}>0$
for sufficiently large $k$ provided that $\delta_{0}<\eta/2$. Now observe that this implies
$2\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}c_{g}(|d|)>k^{-\delta}\end{subarray}}\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)$
$=\left|\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}|c_{g}(|d|)|>k^{-\delta}\end{subarray}}\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)\,+\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}|c_{g}(|d|)|>k^{-\delta}\end{subarray}}\sqrt{\alpha_{g}}c_{g}(|d|)M_{g}(d)\right|$
@@
$2\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}c_{g}(|d|)<-k^{-\delta}\end{subarray}}\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)$
$=\left|\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}|c_{g}(|d|)|>k^{-\delta}\end{subarray}}\sqrt{\alpha_{g}}|c_{g}(|d|)|M_{g}(d)\,-\mathop{{\sum}^{\raisebox{-3.0pt}{\makebox[0.0pt][l]{$\flat$}}}}_{\begin{subarray}{c}x\leq(-1)^{k}d\leq x+H\\ \sqrt{\alpha_{g}}|c_{g}(|d|)|>k^{-\delta}\end{subarray}}\sqrt{\alpha_{g}}c_{g}(|d|)M_{g}(d)\right|$
@@
Thus we have shown that for $\geq(C_{10}-C_{12}-2C_{3}/C_{2}C_{1}^{2})K^{2}$ of the forms $g\in\mathcal{S}_{K}$ the short interval $[x,x+H]$, $x\sim X$, contains numbers $(-1)^{k}d_{\pm}$, with both $d_{\pm}$ odd fundamental discriminants, for which $\sqrt{\alpha_{g}}c_{g}(|d_{+}|)>k^{-\delta}$ and $\sqrt{\alpha_{g}}c_{g}(|d_{-}|)<-k^{-\delta}$, for $\geq(C_{11}-1-C_{2})X$ of the numbers $x\sim X$. Thus we deduce a sign change of $c_{g}(|d|)$ over a positive proportion of intervals of constant size $(C_{1}/C_{7})^{2}$ for a positive proportion of forms $g\in S_{K}$. This leads to